Value-Stream Costing Objective

During the week of June 12, Harrison Manufacturing produced and shipped 16,600 units of its aluminum wheels: 4,400 units of Model A and 12,200 units of Model B. The following costs were incurred:

Materials        Salaries/
Wages
       Machining        Other        Total Cost
Order processing        $18,000                      $18,000
Production planning        64,200                      64,200
Purchasing        27,000                      27,000
Stamping $370,000        37,000        $36,800        $18,400        462,200
Welding 150,000        42,000        42,000        12,000        246,000
Cladding 210,000                             210,000
Testing        10,000                      10,000
Packaging and shipping        8,000                      8,000
Invoicing        14,000                      14,000
  Total $730,000        $220,200        $78,800        $30,400        $1,059,400

Required:

1. Assume initially that the value-stream costs and total units shipped apply only to one model (a single-product value stream). Calculate the unit cost. Round your answer to the nearest dollar amount.
$ per unit

2. Calculate the unit cost for Models A and B. Round your answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $

3. What if Model A is responsible for 40 percent of the materials cost? Calculate the unit materials cost for Models A and B. Round your answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $

Calculate the total unit cost for Models A and B. Round your interim calculations and final answers to the nearest dollar amount.

Unit Cost
Model A $
Model B $
Answer #1
Part- A & B
Unit Cost := Total Cost / Total Unit
=1059400/(4400+12200)= $63.82
Eac Model A & will be the same cost $63.82
Part-C
Model A ( Unit Material Cost )= Total Cost/ Total Unit  
(730000*40%)/4400= $66.36
Model B ( Unit Material Cost )= Total Cost/ Total Unit  
(730000*60%)/12200= $35.90
Computation of Conversion Cost
Unit Conversion Cost =Total Cost/ Total Unit
=(1059400-730000)/(4400+12200)= $19.84
Part-d Computation of Total Unit cost
Particular Material Cost Conversion Cost Total
Model A $66 $20 $86
Model B $36 $20 $56